OpenStudy (anonymous):

Make [2(cos 15 + i sin 15)]^3 to the standard form a+bi

OpenStudy (freckles):

\[[r (\cos(\theta)+i \sin(\theta)]^n \\ =r^n (\cos(n \cdot \theta)+i \sin(n \cdot \theta))\]

OpenStudy (freckles):

\[\text{ distribute } \\ =r^n \cos(n \theta)+ i r^n \sin(n \theta)\]

OpenStudy (freckles):

its called demoive's theorem I think